OpenStudy (anonymous):

x 3 + 6x 2 - 4x - 24

OpenStudy (anonymous):

I want to believe that the equation you're giving has 3x but not x3. therefore if so follow the following steps: 6x2=12; 3x+4x=7x; then carry out subtraction o the following(+12-24)= -12; therefore 7x=12; x=12/7; x=1/1/7

OpenStudy (anonymous):

That is not what my computer showed. sorry

OpenStudy (gorv):

x3+6x2-4x-24=0 x2(x+6)-4(x+6)=0 (x+6)(x2-4)=0 x+6=0 & x2-4=0 x=-6 & x2=4 implies x = 2 & x= -2